# NeuroNAS: A Framework for Energy-Efficient Neuromorphic Compute-in-Memory Systems using Hardware-Aware Spiking Neural Architecture Search

### v2

@@ V-C Searching Time Speed-Up
Fig. 13 presents the experimental results of the searching time. In general, our NeuroNAS offers significant speed-ups for the searching time as compared to the state-of-the-art. Specifically, NeuroNAS improves the searching time by 4.8x-8.4x for CIFAR-10, by 3.1x-7.2x for CIFAR-100, and by 4.4x-8.9x for TinyImageNet-200. The reason is that, our NeuroNAS employs a fewer SNN operations, minimizes redundant architecture candidates to evaluate, incorporates multiple constraints into the search process, hence leading to a smaller search space and a faster searching time as compared to the state-of-the-art.
In summary, our NeuroNAS can quickly provide an SNN that fulfills the given memory, area, latency, and energy constraints. For instance, NeuroNAS_FxP8 maintains high accuracy comparable to the state-of-the-art, while achieving up to 92% area savings, 1.2x latency speed-ups, 84% energy savings, and 6.6x search time speed-ups.
@@ VI Conclusion
We propose the NeuroNAS framework to find an SNN architecture that achieves high accuracy under multiple constraints. It is achieved by optimizing SNN operations, developing an efficient SNN using novel QaFE policy, and designing a hardware-aware search algorithm that includes the given constraints in its NAS. As a result, NeuroNAS maintains high accuracy comparable to state-of-the-art while improving other aspects to meet all constraints. For instance, NeuroNAS_FxP8 achieves up to 92% area savings, 1.2x latency speed-ups, 84% energy savings, and 6.6x search time speed-ups across all datasets. In this manner, our NeuroNAS framework enables the efficient design automation for developing neuromorphic CIM systems for diverse application use-cases.

### v3

@@ V-C Searching Time Speed-Up
Fig. 13 presents the experimental results of the searching time. In general, our NeuroNAS offers significant speed-ups for the searching time as compared to the state-of-the-art. Specifically, NeuroNAS improves the searching time by 4.8x-8.4x for CIFAR-10, by 3.1x-7.2x for CIFAR-100, and by 4.4x-8.9x for TinyImageNet-200; see 6. The reason is that, our NeuroNAS employs a fewer SNN operations, minimizes redundant architecture candidates to evaluate, incorporates multiple constraints into the search process, hence leading to a smaller search space and a faster searching time as compared to the state-of-the-art.
V-D Comparison to Conventional ANN-based Solutions
In ANN domain, the prominent state-of-the-art work in “NAS without training” that considers the same benchmark with neural cell-based strategy (i.e., NAS-Bench-201 [25]) is NASWOT [24]. We observe that NASWOT with a 32-bit precision can achieve around 92.81% accuracy for CIFAR-10 and around 69.35% accuracy for CIFAR-100. Meanwhile, NeuroNAS with a 32-bit precision achieves around 91.58% for CIFAR-10 and around 64.04% accuracy for CIFAR-100. These results indicate that, our NeuroNAS can obtain competitive accuracy as compared to the conventional ANNs because of its effective searching strategy, thus providing an SNN architecture that can learn well during training process.
Besides competitive accuracy, NeuroNAS also offers significantly higher energy efficiency than ANN-based solutions due to the nature of event-based SNN processing as well as the reduced off-chip data movements, which are maximized through the employment of IMC hardware; see Table II and Fig. 14. Specifically, our NeuroNAS_FxP16 can achieve 86% energy saving while maintaining accuracy within 2% from APNAS96 [32] for CIFAR-10 dataset.
TABLE II: Comparison to ANN-based solutions from APNAS [32] for network size, accuracy, and energy consumption considering the CIFAR-10 dataset and 65nm technology nodes.
Domain
NAS Method
Number of
Parameters
Accuracy
Energy
Consumption
[…]
Figure 14: Trade-offs among accuracy, network size (number of parameters), and energy consumption for ANN-based solutions and our NeuroNAS.
@@ VI Conclusion
We propose the novel NeuroNAS framework to develop an SNN architecture for a neuromorphic IMC platfom that can achieve high accuracy under multiple requirements and constraints. Its key steps include: (1) optimizing SNN operations, (2) developing an efficient SNN using novel QaFE policy, and (3) employing a hardware-aware search algorithm that incorporates all constraints in its NAS process. As a result, NeuroNAS maintains high accuracy comparable to the state-of-the-art while improving other aspects to fulfill all given constraints. For instance, NeuroNAS_FxP8 achieves up to 92% area savings, 1.2x latency improvements, 84% energy savings, and 6.6x search time speed-ups across all datasets. Therefore, NeuroNAS framework enables an effective design automation for streamlining the development of energy-efficient neuromorphic IMC for intelligent mobile agents.